# Similarity of Neural Architectures using Adversarial Attack Transferability

### v3

@@ Abstract
In recent years, many deep neural architectures have been developed for image classification. Whether they are similar or dissimilar and what factors contribute to their (dis)similarities remains curious. To address this question, we aim to design a quantitative and scalable similarity measure between neural architectures. We propose Similarity by Attack Transferability (SAT) from the observation that adversarial attack transferability contains information related to input gradients and decision boundaries widely used to understand model behaviors. We conduct a large-scale analysis on 69 state-of-the-art ImageNet classifiers using our proposed similarity function to answer the question. Moreover, we observe neural architecture-related phenomena using model similarity that model diversity can lead to better performance on model ensembles and knowledge distillation under specific conditions. Our results provide insights into why developing diverse neural architectures with distinct components is necessary.
@@ 1 Introduction
The advances in the deep neural networks (DNN) architecture design have taken a key role in their success by making the learning process easier (e.g., normalization [50, 109, 2] or skip connection [40]), enforcing human inductive bias in the architecture [57], or increasing model capability with the self-attention mechanism [101]. With different architectural components containing architectural design principles and elements, a number of different neural architectures have been proposed. They have different accuracies, but several researches have pointed out that their predictions are not significantly different [69, 68, 33].
@@
By this, can we say that recently developed DNN models with different architectural components are similar or the same? The answer is no. It is because a model prediction is not the only characteristic to compare their similarities. Existing studies have found differences by focusing on different features, such as layer-by-layer network component [55, 77], a high-level understanding by visualization of loss surface [27], input gradient [86, 84], and decision boundary [85]. Researchers could understand the similarity between models through these trials; however, the similarity comparison methods from previous studies are insufficient for facilitating comprehensive studies because they do not satisfy two criteria that practical metrics should meet: (1) providing a quantitative similarity score and (2) being compatible with different base architectures (e.g., CNN and Transformer). Recently, Somepalli et al. [85] suggested a quantitative similarity metric based on the on-manifold decision boundary. However, this method needs heavy computation due to the triplet sampling process when the number of classes becomes large (e.g., ImageNet dataset with 1,000 classes). If we limit the total computations, this method will suffer from inaccurate approximation with high variance.
Figure 1: t-SNE plot showing 10 clusters of 69 neural networks using our similarity function, SAT. We cluster 69 neural networks in 10 groups based on SAT. The group number follows the result of spectral clustering (Tab. 2). Due to the space limit, we only illustrate representative models in each group.
@@
This paper proposes a quantitative similarity function that is scalable and easily applicable to diverse architectures, named Similarity by Attack Transferability (SAT). We focus on adversarial attack transferability (AT), which indicates how generated adversarial perturbation is transferable between two different architectures. Specifically, it is widely studied that the vulnerable points of neural architectures depend on their own architectural property or how models capture the features from inputs, such as the usage of self-attention [31], the stem layer [48], and the model dependency on high or low-frequency components of input [3, 54]. Thus, if two different neural architectures are similar, the AT between the models is high because they share similar vulnerability [79]. Furthermore, AT can be a reliable approximation for comparing the input gradients [67], decision boundary [53], and loss landscape [25] of models. All of them are widely-used frameworks to understand model behavior and the difference between models and used to measure the similarity of models in previous works [90, 83, 86, 84, 89, 4, 16, 85, 27, 61, 73]; in other words, we can capture various model characteristics by SAT.
@@
We quantitatively measure pairwise SATs of 69 different ImageNet-trained neural architectures, provided by Wightman [107]. With SAT similarity scores of 69 architectures, we analyze what components among 13 architectural components (e.g., normalization, activation, the design choice for stem layers) that consist of neural architectures largely affect model diversity by gradient boosting regression and clustering on model similarities. Furthermore, we observe relationships between model similarity and model ensemble performances, and distillation performances. The main contributions of this work can be summarized as follows.
  • We propose a scalable and quantitative similarity between neural architectures, named Similarity by Attack Transferability (SAT). SAT is easy to conduct and reports stable results even when there is a change in data sampling.
  • We conduct a comprehensive analysis using 69 models to investigate what components make architectures behave similarly or differently. To the best of our knowledge, this is the first large-scale analysis exploring the relationship between model similarity and architectural components. Furthermore, our findings are not limited to architectural components but also encompass training strategies; different training strategies do not affect model diversity as much as architectural differences.
  • We identify several outgrowths using SAT. First, we find model dissimilarity leads to better ensemble performance and analyze the reasons behind this. Additionally, we suggest new guidelines for selecting a teacher model suited to the student model in knowledge distillation. These observations provide insights into why we need to develop diverse architectures.
@@ 2 Related work
Similarity between DNNs has been actively explored recently. Several studies focused on comparing intermediate features for understanding the behavior of DNNs. Raghu et al. [77] observed the difference between layers, training methods, and architectures (e.g., CNN and ViT) based on layer-by-layer comparison [55]. Some studies have focused on loss landscapes, which helps to visualize the loss of models on the parameter space [27, 61, 73]. Although these methods show a visual inspection, they cannot support quantitative measurements. In this paper, we aim to support a quantitative similarity function.
Another line of research has been focused on prediction-based statistics, e.g., comparing wrong and correct predictions [58, 32, 33, 81]. However, as recent complex DNNs are getting almost perfect, just focusing on prediction values can be misleading. For example, Meding et al. [69] observed that recent DNNs show highly similar predictions. In this case, prediction-based methods will be no more informative. On the other hand, our SAT can provide meaningful information and findings for 69 recent neural architectures.
Input gradient is another popular framework to understand model behavior by observing how a model will change predictions by local pixel changes [83, 86, 84, 89, 4]. If two models are similar, their input gradients will also be similar. These methods are computationally efficient, and no additional training is required; they can provide a visual understanding of the given input. However, input gradients are inherently noisy; thus, these methods will need additional pre-processing, such as smoothing, for a stable computation [16]. Also, input gradient-based quantity measures usually measure how the input gradient matches the actual foreground, i.e., we need ground-truth foreground masks for measuring such scores. On the contrary, SAT needs no additional pre-processing and mask annotations.
Comparing the decision boundaries of models will provide a high-level understanding of how models behave differently for input changes and how models extract features from complicated data dimensions. Somepalli et al. [85] analyzed characteristics of models by comparing their decision boundaries on the on-manifold plane constructed by three random images. However, we observe that this approach derives unstable measurements on huge datasets (e.g., ImageNet) and is expensive compared to SAT.
@@ 3 Similarity by Attack Transferability (SAT)
In this section, we propose a quantitative similarity measure between two neural architectures using adversarial attack transferability, which indicates whether an adversarial sample from a model can fool another model or not. The concept of adversarial attack has effectively pointed out the vulnerabilities of DNNs by input gradient [91, 34, 67].
Interestingly, these vulnerabilities have been observed to be intricately linked to architectural properties. Fu et al. [31] demonstrated the effect of the attention modules in architecture in terms of attack success rate. Hwang et al. [48] analyzed that the stem layer structure of models causes models to have different adversarial vulnerable points in the input space, e.g., video action-recognition models periodically have vulnerable frames, such as every four frames. In other words, an adversarial sample to a model highly depends on the inherent architectural property of the model, e.g., the base architecture or the stem layer structure.
Another perspective emphasized the dissimilarities in dependencies on high-frequency and low-frequency components between CNN-based and transformer-based models, showing different vulnerabilities to different adversarial attacks [3, 54]. As shown by Park and Kim [73], different architectural choices behave as different frequency filters (e.g., the self-attention mechanism works as a low-pass filter, while the convolution operation works as a high-pass filter); thus, we can expect that the different architectural component choice will affect to the model vulnerability, e.g., vulnerability to high-frequency perturbations. In other words, if we can measure how the adversarial vulnerabilities of the models are different, we also can measure how the networks are similar or dissimilar.
To measure how model vulnerabilities differ, we employ adversarial transferability, where it indicates whether an adversarial sample from a model can fool another model. If two models are more similar, their adversarial transferability gets higher [79, 25, 62]. On the other hand, because the adversarial attack targets vulnerable points varying by architectural components of DNNs [47, 31, 48, 54], if two different models are dissimilar, the adversarial transferability between them gets lower. Furthermore, attack transferability can be a good approximation for measuring the differences in input gradients [67], decision boundaries [53], and loss landscape [25], where they are widely used techniques for understanding model behavior and similarity between models as discussed in the related work section. While previous approaches are limited to non-quantitative analysis, inherent noisy property, and computational costs, adversarial transferability can provide quantitative measures with low variances and low computational costs.
@@
$\text{SAT}(A,B)=\log\big{[}\max\big{\{}\varepsilon_{s},100\times$ (1)
$\frac{1}{2|X_{AB}|}\sum_{x\in X_{AB}}\left\{\mathbb{I}({A(x_{B})}\neq y)+\mathbb{I}({B(x_{A})}\neq y)\right\}\big{\}}\big{]},$
@@
where $\varepsilon_{s}$ is a small scalar value to prevent the ill-posedness of $\text{SAT}(\cdot)$. If $A=B$ and we have an oracle adversary, then $\text{SAT}(A,A)=\log 100$. In practice, a strong adversary (e.g., PGD [67] or AutoAttack [22]) can easily achieve a nearly-zero accuracy if a model is not trained by an adversarial attack-aware strategy [67, 21]. On the other hand, if the adversarial attacks on $A$ are not transferable to $B$ and vice versa, then $\text{SAT}(A,B)=\log\varepsilon_{s}$.
@@
SAT (Eq. 1) can act as a similarity function between models. Ideally, we would like to define a similarity function $d$ between two models with the following three properties: (1) $n=\arg\min_{m}d(n,m)$, (2) $d(n,m)=d(m,n)$ and (3) $d(n,m)>d(n,n)$ if $n\neq m$. The proposed function satisfies these properties if we assume the perfect adversary. If the adversary is perfect, then $\text{acc}_{A\rightarrow A}$ will be zero, and it will be the minimum because accuracy is non-negative. “$\text{acc}_{A\rightarrow B}+\text{acc}_{B\rightarrow A}$” is symmetric thereby SAT is symmetric. Finally, SAT will satisfy $d(n,m)\geq d(n,n)$ if $n\neq m$ where it is a weaker condition than (3).
Now, we extensively analyze 69 recent state-of-the-art ImageNet-trained networks using our SAT score by focusing on two questions. (1) Which network component contributes to the diversity between models? (2) Why do we need to develop various neural architectures?
@@
Settings. We use the PGD attack for the adversary. We set the iteration to 50, the learning rate to 0.1, and $\varepsilon$ to $8/255$. We select 69 neural architectures trained on ImageNet [80] from the PyTorch Image Models library [107]. To reduce the unexpected effect of a significant accuracy gap, the chosen model candidates are limited to the models whose top-1 accuracy is between 79% and 83%. We also ignore the models with unusual training techniques, such as training on extra training datasets, using a small or large input resolution (e.g., less than 200 or larger than 300), or knowledge distillation. When $A$ and $B$ take different input resolutions, then we resize the attacked image from the source network for the target network. We also sub-sample 10% ImageNet validation images (i.e., 5,000 images) to measure the similarity. This strategy makes our similarity score more computationally efficient. The full pairwise similarities of 69 models can be found in Appendix.
Ablation study on attack methods. In our analyses, we compute SAT using the famous PGD attack. To explore the effect of the choice of attack methods, we compute SAT using two different gradient-based attack methods; Autoattack [22], the state-of-the-art attack method based on the ensemble of four diverse attacks, and Patchfool attack [31], specially designed for vision transformers. We sample 8 representative models among 69 models for testing the effect of Autoattack on SAT: ViT-S, CoaT-Lite Small, ResNet-101, LamHaloBotNet50, ReXNet-150, NFNet-L0, Swin-T and Twins-pcpvt. Fig. 1(a) shows the high correlation between SAT scores using PGD and Autoattack; it shows a correlation coefficient of 0.98 with a p-value of $1.43\times 10^{-18}$. For testing the Patchfool attack, we only generate adversarial perturbations on ViT-S and get attack transferability to all other models (68 models) because it only targets Transformers. In Fig. 1(b), Patchfool also shows a high correlation compared to the PGD attack (correlation coefficient 0.91 with p-value $3.62\times 10^{-27}$). The results show that SAT score is robust to the choice of attack methods if the attack is strong enough. Note that PatchFool needs a heavy modification on the model code to extract attention layer outputs manually. Contrarily, PGD and AutoAttack are model-agnostic and need no model code modification. Thus, as PatchFool and PGD show similar rankings, we use PGD due to its simplicity.
(a) PGD vs. Autoattack
(b) PGD vs. Patchfool
Figure 2: Effect of different adversarial attack methods to SAT. The trend line and its 90% confidence interval are shown. We show the relationship between our SAT using PGD and SAT using other attacks, (a) Autoattack and (b) Patchfool attack.
Comparison with other similarities. Somepalli et al. [85] proposed a sampling-based similarity score for comparing decision boundaries of models. SAT has two advantages over Somepalli et al.: computational efficiency and the reliability of the results. First, SAT involves sampling 5,000 images and using 50-step PGD; the computation cost is [5,000 (sampled images) $\times$ 50 (PGD steps) + 5,000 (test to the other model)] $\times$ 2 (two models). On the other hand, Somepalli et al. [85] sample 500 triplets and generate 2,500 points to construct decision boundaries. In this case, the whole amount of inference is [500 (sampled triplets) $\times$ 2,500 (grid points)] $\times$ 2 (two models), which is 4.9 times larger than our method. Secondly, Somepalli et al. [85] sampled three images of different classes. As the original paper used the CIFAR-10 dataset [56], 500 triplets can cover all possible combinations of three classes among the ten classes (($10\choose 3$ = 120 $<$ 500). However, in datasets with many classes, it becomes computationally infeasible to represent all possible combinations of three classes (e.g., ImageNet needs ($1000\choose 3$ = 164,335,500, almost 130 times greater than the number of training examples of ImageNet). Additionally, we find that the similarity results are unreliable (i.e., showing a high variance) when the number of sample triplets is small. In Tab. 1, we calculate the similarity scores between ConvNeXt-T [66] and Swin-T [65] from ten different sets with varying sample sizes. SAT exhibits significantly better stability (i.e., low variances) than Somepalli et al. Note that we use our similarity measurement as the percentage degree without the logarithmic function and control the scale of samples to maintain similar computation complexity between SAT and the compared method for a fair comparison.
Somepalli et al. [85] SAT (ours)
# triplets std cost # images std cost
10 (4.8×$\times 10^{-13}$) 4.49 1.0 500 (0.01) 1.88 1.0
20 (9.6×$\times 10^{-13}$) 3.28 2.0 1000 (0.02) 1.05 2.0
50 (2.4×$\times 10^{-12}$) 1.63 5.0 2500 (0.05) 0.91 5.2
100 (4.8×$\times 10^{-12}$) 1.54 10.0 5000 (0.1) 0.77 10.2
Table 1: Comparison of stability of measurements. We compare the stability of method by [85] and our method. Stability is indicated by the standard deviation (std). The numbers in ($\cdot$) mean the sampling ratio to all possible combinations to compute the exact value. “cost” denotes relative costs compared to the total forward costs for the 50K ImageNet validation set: Somepalli et al. needs ($50\text{K}\choose 3$ = 2.1×$\times 10^{13}$ and SAT needs 50K. Here, we assume that forward and backward computations cost the same.

### v4

@@ Abstract
In recent years, many deep neural architectures have been developed for image classification. Whether they are similar or dissimilar and what factors contribute to their (dis)similarities remains curious. To address this question, we aim to design a quantitative and scalable similarity measure between neural architectures. We propose Similarity by Attack Transferability (SAT) from the observation that adversarial attack transferability contains information related to input gradients and decision boundaries widely used to understand model behaviors. We conduct a large-scale analysis on 69 state-of-the-art ImageNet classifiers using our SAT to answer the question. In addition, we provide interesting insights into ML applications using multiple models, such as model ensemble and knowledge distillation. Our results show that using diverse neural architectures with distinct components can benefit such scenarios.
Keywords:
Architecture Similarity Adversarial Attack Transferability
Figure 1: t-SNE plot showing 10 clusters of 69 neural networks using our similarity function, SAT.
@@ 1 Introduction
The advances in deep neural networks (DNN) architecture design have taken a key role in their success by making the learning process easier (e.g., normalization [52, 116, 3] or skip connection [42]), enforcing human inductive bias [60], or increasing model capability with the self-attention mechanism [106]. With different architectural components containing architectural design principles and elements, a number of different neural architectures have been proposed. They have different accuracies, but several researches have pointed out that their predictions are not significantly different [72, 71, 35].
@@
By this, can we say that recently developed DNN models with different architectural components are similar or the same? The answer is no. It is because a model prediction is not the only characteristic to compare their similarities. Existing studies have found differences by focusing on different features, such as layer-by-layer network component [58, 82], a high-level understanding by visualization of loss surface [28], input gradient [91, 89], and decision boundary [90]. Researchers could understand the similarity between models through these trials; however, the similarity comparison methods from previous studies are insufficient for facilitating comprehensive studies because they do not satisfy two criteria that practical metrics should meet: (1) providing a quantitative similarity score and (2) being compatible with different base architectures (e.g., CNN and Transformer). Recently, Tramèr et al. [103] and Somepalli et al. [90] suggested a quantitative similarity metric based on measuring differences in decision boundaries. However, these methods have limitations due to the non-tractable decision boundaries and limited computations as shown in Sec. 3.
@@
We propose a quantitative similarity that is scalable and easily applicable to diverse architectures, named Similarity by Attack Transferability (SAT). We focus on adversarial attack transferability (AT), which indicates how generated adversarial perturbation is transferable between two different architectures. It is widely studied that the vulnerability of DNNs depends on their own architectural property or how models capture the features from inputs, such as the usage of self-attention [33], the stem layer [50], and the dependency on high or low-frequency components of input [4, 57]. Thus, if two different models are similar, the AT between the models is high because they share similar vulnerability [84]. Furthermore, AT can be a reliable approximation for comparing the input gradients [70], decision boundary [56], and loss landscape [26]. All of them are widely-used frameworks to understand model behavior and differences between models and used to measure the similarity of models in previous works [91, 89, 94, 6, 18, 90, 28, 64, 103, 113]; namely, SAT can capture various model properties.
@@
We quantitatively measure pairwise SATs of 69 different ImageNet-trained neural architectures from [114]. We analyze what components among 13 architectural components (e.g., normalization, activation, …) that consist of neural architectures largely affect model diversity. Furthermore, we observe relationships between SAT and practical applications, such as ensemble and distillation.
@@ 2 Related Work
Similarity between DNNs has been actively explored recently. Several studies focused on comparing intermediate features to understand the behavior of DNNs. Raghu et al. [82] observed the difference between layers, training methods, and architectures (e.g., CNN and ViT) based on layer-by-layer comparison [58]. Some studies have focused on loss landscapes by visualizing the loss of models on the parameter space [28, 64, 78]. Although these methods show a visual inspection, they cannot support quantitative measurements. On the other hand, our goal is to support a quantitative similarity by SAT.
Another line of research has been focused on prediction-based statistics, e.g., comparing wrong and correct predictions [61, 34, 35, 86]. However, as recent complex DNNs are getting almost perfect, just focusing on prediction values can be misleading; Meding et al. [72] observed that recent DNNs show highly similar predictions. In this case, prediction-based methods will be no more informative. Meanwhile, our SAT can provide meaningful findings for 69 recent NNs.
Input gradient is another popular framework to understand model behavior by observing how a model will change predictions by local pixel changes [88, 91, 89, 94, 6]. If two models are similar, their input gradients will also be similar. These methods are computationally efficient, and no additional training is required; they can provide a visual understanding of the given input. However, input gradients are inherently noisy; thus, these methods will need additional pre-processing, such as smoothing, for a stable computation [18]. Also, these methods usually measure how the input gradient matches the actual foreground, i.e., we need ground-truth foreground masks for measuring such scores. On the contrary, SAT needs no additional pre-processing and mask annotations.
Comparing the decision boundaries will provide a high-level understanding of how models behave differently for input changes and how models extract features from complicated data dimensions. Recent works [103, 113] suggested measuring similarity by comparing distances between predictions and decision boundaries. Meanwhile, Somepalli et al. [90] analyzed models by comparing their decision boundaries on the on-manifold plane constructed by three random images. However, these approaches suffer from inaccurate approximation, non-tractable decision boundaries, and finite pairs of inputs and predictions.
Finally, different behaviors of CNNs and Transformers have been studied in specific tasks, such as robustness [5, 74], layer-by-layer comparison [82, 78] or decision-making process [53]. Our work aims to quantify the similarity between general NNs, not only focusing on limited groups of architecture.
@@ 3 Similarity by Attack Transferability (SAT)
Here, we propose a quantitative similarity between two architectures using adversarial attack transferability, which indicates whether an adversarial sample from a model can fool another model. The concept of adversarial attack has effectively pointed out the vulnerabilities of DNNs by input gradient [95, 36, 70].
Interestingly, these vulnerabilities have been observed to be intricately linked to architectural properties. For example, Fu et al. [33] demonstrated the effect of the attention modules in architecture on attack success rate. Hwang et al. [50] analyzed that the stem layer structure causes models to have different adversarial vulnerable points in the input space, e.g., video models periodically have vulnerable frames, such as every four frames. Namely, an adversarial sample to a model highly depends on the inherent architectural property of the model.
Another perspective emphasized the dissimilarities in dependencies on high-frequency and low-frequency components between CNN-based and transformer-based models, showing different vulnerabilities to different adversarial attacks [4, 57]. Different architectural choices behave as different frequency filters (e.g., the self-attention works as a low-pass filter, while the convolution works as a high-pass filter) [78]; thus, we can expect that the different architectural component choices will affect the model vulnerability, e.g., vulnerability to high-frequency perturbations. If we can measure how the adversarial vulnerabilities of the models are different, we also can measure how the networks are dissimilar.
To measure how model vulnerabilities differ, we employ adversarial attack transferability (AT), where it indicates whether an adversarial sample from a model can fool another model. If two models are more similar, their AT gets higher [84, 26, 65]. On the other hand, because the adversarial attack targets vulnerable points varying by architectural components of DNNs [49, 33, 50, 57], if two different models are dissimilar, the AT between them gets lower. Furthermore, attack transferability can be a good approximation for measuring the differences in input gradients [70], decision boundaries [56], and loss landscape [26], where they are widely used techniques for understanding model behavior and similarity between models as discussed in the related work section. While previous approaches are limited to non-quantitative analysis, inherent noisy property, and computational costs, adversarial transferability can provide quantitative measures with low variances and low computational costs.
@@
$\text{SAT}(A,B)=\log\big{[}\max\big{\{}\varepsilon_{s},100\times\frac{1}{2|X_{AB}|}\sum_{x\in X_{AB}}\left\{\mathbb{I}({A(x_{B})}\neq y)+\mathbb{I}({B(x_{A})}\neq y)\right\}\big{\}}\big{]},$ (1)
@@
where $\varepsilon_{s}$ is a small scalar value. If $A=B$ and we have an oracle adversary, then $\text{SAT}(A,A)=\log 100$. In practice, a strong adversary (e.g., PGD [70] or AutoAttack [23]) can easily achieve a nearly-zero accuracy if a model is not trained by an adversarial attack-aware strategy [70, 22]. Meanwhile, if the adversarial attacks on $A$ are not transferable to $B$ and vice versa, then $\text{SAT}(A,B)=\log\varepsilon_{s}$.
@@
Ideally, we aim to define a similarity $d$ between two models with the following properties: (1) $n=\arg\min_{m}d(n,m)$, (2) $d(n,m)=d(m,n)$ and (3) $d(n,m)>d(n,n)$ if $n\neq m$. If the adversary is perfect, then $\text{acc}_{A\rightarrow A}$ will be zero, and it will be the minimum because accuracy is non-negative. “$\text{acc}_{A\rightarrow B}+\text{acc}_{B\rightarrow A}$” is symmetric thereby SAT is symmetric. Finally, SAT satisfies $d(n,m)\geq d(n,n)$ if $n\neq m$ where it is a weaker condition than (3).
Comparison with other methods.
[…]
In this scenario, we can choose three strategies. First, we can count the number of samples whose predicted labels are different for given $x$, which is prediction-based measurements or Somepalli et al. [90]. As we assumed sparsity of $x$, this approach cannot measure the area of uncovered $x$ domain, hence, its approximation will be incorrect (purple box in Fig. 2) or needs too many perturbations to search uncovered $x$. In Sec. 0.A.1, we empirically show that Somepalli et al. [90] suffers from the high variance even with a large number of samples while SAT shows a low variance with a small number of samples.
Figure 2: How SAT works? Conceptual figure to understand SAT by the lens of the decision boundary. Each line denotes the decision boundary of a binary classification model, and each dot denotes individual prediction for given inputs.
Second, we can measure the minimum distance between $f_{p}(x)$ and $f_{d}$ as Tramèr et al. [103]. This only measures the distance to its closest decision boundary without considering the other model. The yellow box of Fig. 2 shows if two predictions are similar at $x$, it would compute an approximation of $|f_{d}(x)-g_{d}(x)|$ for $x$. However, if two predictions are different, it will compute a wrong approximation. Moreover, in practice, searching $\epsilon$ is unstable and expensive.
Lastly, we can count the number of non-transferred adversarial samples (red box in Fig. 2), which is our method, SAT. If we have an oracle attack method that exactly moves the point right beyond the decision boundary, our SAT will measure the $\ell_{0}$ approximation of $\min(|f_{d}(x)-g_{d}(x)|,\epsilon)$ for given $x$. Namely, SAT can measure whether two decision boundaries are different by more than $\epsilon$ for each $x$. If we assume that the difference between decision boundaries is not significantly large and $\epsilon$ is properly chosen, SAT will compute an approximated decision boundary difference. We also compare SAT and other methods from the viewpoint of stability and practical usability in Sec. 5.1 and Appendix 0.A.
Discussions.
In practice, we do not have an oracle attack method. Instead, we employ the PGD attack [70] as the adversarial attack method. In Sec. 0.B.1, we investigate the robustness of SAT to the choice of the attack methods. In summary, SAT measured by PGD shows a high correlation with SAT measured by various attacks, e.g., AutoAttack [23], attacks designed for enhancing attack transferability, such as MIFGSM [29] and VMIFGSM [112], low-frequency targeted attacks, such as low-frequency PGD [38], method-specific attacks, such as PatchPool [33], or generative model-based attacks, such as BIA [129].
Also, SAT assumes an optimal attack with proper $\epsilon$. However, this assumption can be broken under the adversarial training setting when we use a practical attacker. Also, as shown by Tsipras et al. [105] and Ilyas et al. [51], adversarial training will lead to a different decision boundary from the original model. In Sec. 0.B.2, we empirically investigate the effect of adversarial training to SAT. We observe that different adversarial training methods make as a difference as different training techniques, which we will discuss in Sec. 4.2.
Analyzing 69 models.
@@
Now, we analyze 69 recent ImageNet classifiers using SAT by focusing on two questions. (1) Which network component contributes to the diversity between models? (2) Why do we need to develop various neural architectures? The full list of the architectures can be found in Appendix 0.C. We use the PGD attack [70] for the adversary. We set the iteration to 50, the learning rate to 0.1, and $\varepsilon$ to $8/255$. As we discussed earlier, we show that SAT is robust to the choice of the adversarial attack method. We select 69 neural architectures trained on ImageNet [85] from the PyTorch Image Models library [114]. To reduce the unexpected effect of a significant accuracy gap, the chosen model candidates are limited to the models whose top-1 accuracy is between 79% and 83%. We also ignore the models with unusual training techniques, such as training on extra training datasets, using a small or large input resolution (e.g., less than 200 or larger than 300), or knowledge distillation. When $A$ and $B$ take different input resolutions, then we resize the attacked image from the source network for the target network. We also sub-sample 10% ImageNet validation images (i.e., 5,000 images) to measure the similarity. This strategy makes our similarity score more computationally efficient.